Repeating decimals(Multiplication and division) work out by steps please
(a.)o.333...x6/7
(b.)0.888...divided by 2/9
)c.) 0.2727...x 8/11
(d.) 0.1212.. divided by 5/3

Answer by Alan3354(69443)

Repeating decimals(Multiplication and division) work out by steps please
(a.)o.333...x6/7
Find the fraction of the repeating decimal first.
n = 0.333...
10n = 3.333...
-1n = 0.333...
-------------------- Subtract
9n = 3
n = 1/3
-----
(1/3)*(6/7) = 3/7
============================
(b.)0.888...divided by 2/9
Similar to (a), you try it.
--------------
(c.) 0.2727...x 8/11
n = 0.2727...
100n = 27.2727...
- 1n = -0.2727...
-----------------
99n = 27
n = 3/11
==============
(d.) 0.1212.. divided by 5/3
=================
Similar to (c)
= 0.1212... times 3/5